OpenStudy (anonymous):

Help will medal and fan!! The angle of elevation from the end of the shadow to the top of the building is 70 degrees and the distance is 180 feet. Find the height of the building to the nearest foot. Find the length of the shadow to the nearest foot.

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

@boricua5

OpenStudy (anonymous):

@Secret-Ninja

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

So we're looking for the opposite of the right triangle.

OpenStudy (secret-ninja):

Looks like this also may have already been answered. It shows the work too. http://openstudy.com/study#/updates/52a26ccbe4b088e695c13de9

OpenStudy (igreen):

We can use tangent.. \(\sf Tangent = \dfrac{Opposite}{Adjacent}\) We are given the angle measurement and the adjacent. We can write: \(\sf tan(70^o) = \dfrac{x}{180}\) Now we find tan 70 degrees and multiply that by 180.

OpenStudy (anonymous):

Thank you both!!